### v2

# Improving Neutral Point-of-View Generation with Data- and Parameter-Efficient RL

Abstract
The paper shows that parameter-efficient reinforcement learning (PE-RL) is a highly effective training regime to improve large language models’ (LLMs) ability to answer queries on sensitive topics with a Neutral Point of View (NPOV), i.e. to provide significantly more informative, diverse and impartial answers. This is shown by evaluating PE-RL and multiple strong baselines—including LoRA finetuning (strongest baseline), SFT and RLHF. PE-RL not only improves on overall NPOV quality compared to the strongest baseline ($97.06\%\rightarrow 99.08\%$), but also scores much higher on features linguists identify as key to separating sufficient answers from “great” answers ($60.25\%\rightarrow 85.21\%$ for presence of supportive details, $68.74\%\rightarrow 91.43\%$ for absence of oversimplification). A qualitative analysis corroborates this. Moreover, our evaluation also finds a key property of PE-RL for this task: unlike methods that update all parameters, it generalises out of topic. Finally, to enable further studies we also release the dataset, SHQ-NPOV, and provide a methodology to create such datasets through iterative rounds of human peer-critique and annotator training.
Improving Neutral Point-of-View Generation with Data- and Parameter-Efficient RL
Jessica Hoffmann1, Christiane Ahlheim2, Zac Yu2, Aria Walfrand2, Jarvis Jin1, Marie Tano2, Ahmad Beirami1, Erin van Liemt2, Nithum Thain1, Hakim Sidahmed1, Lucas Dixon1 1Google DeepMind, 2Google Correspondence: jhoffmann@google.com
1 Introduction
Figure 1: Pipeline to create Neutral Point of View (NPOV) answers to queries on sensitive topics.
A long standing goal of AI question-answering systems has been to provide multiple perspectives to controversial queries Cardie et al. (2003); Ku and Chen (2007); Sun et al. (2017); Chen et al. (2022, 2019). Applications aimed at summarizing diverse opinions range from helping people make more informed choices on product purchases Hu and Liu (2004) to predicting electionsTumasjan et al. (2010). With the development of chatbots and LLMs playing an increasingly significant role in search and question answering, several authors such as Metzler et al. (2021) have argued that there is growing importance to provide diverse, unbiased points of view. Doing this well may help with the challenges of polarization and misinformation Costello et al. (2024), and failure has the risk of exacerbating social tensions. This motivates the question of how to efficiently and effectively control LLMs to provide high quality, multi-perspective responses to controversial topics.
In this paper, we addresses this challenge by first describing a methodology we developed to create a small high quality dataset of responses to controversial topics with varying degrees of neutrality; and second by investigating training regimes to identify one that not only learns to generate high quality, neutral, multi-perspective responses, but also successfully generalises to out of domain topics. The resulting dataset we use as a case study for our method consists of 300 examples with high inter-annotator agreement, and is based on adapting Wikipedia’s Neutral Point of View (NPOV) policy. It includes a diverse spectrum of responses exhibiting varying degrees of bias, informativeness, clarity of the arguments, and over-simplification, among other features.
The key contribution is the identification of parameter-efficient reinforcement learning (PE-RL) Sidahmed et al. (2024) as a remarkable effective, and understudied paradigm for low-data controlled generation. PE-RL both improves on overall NPOV quality compared to the strongest baseline ($97.06\%\rightarrow 99.08\%$), and also scores much higher on features which linguists identify as being key to separating sufficient answers from great answers ($60.25\%\rightarrow 85.21\%$), for presence of supportive details ($68.74\%\rightarrow 91.43\%$), and for absence of oversimplification. We compare this methods to strong baselines, including LoRA SFT (parameter-efficient Supervised FineTuning), best-of-10, traditional RLHF (Reinforcement Learning from Human Feedback) and traditional SFT (Supervised FineTuning). Perhaps the most surprising result of our evaluation is that PE-RL, unlike the non-parameter-efficient methods (including RLHF), exhibits no perceptible over-fitting, despite the small-data regime during training. When evaluated on 300 additional questions spanning 30 different topics sourced from a different set of sites, the PE-RL based LLM generator exhibits equally good responses.
While our work makes significant advances in the feasibility of using LLMs to provide multiple nuanced-perspectives to help people understand the diversity of human views, it does not address a number of important and hard open questions that would be necessary for broader application of these methods. Notably, we do not address how to source the perspectives, what the right definition of neutrality should be, and whether our adaptation of Wikipedia’s NPOV is the best way to align AI chatbots or search responses to broader societal benefit.
Part of the challenge with making progress on these open research questions has been the need for large and expensive-to-develop datasets to tune LLMs for a given formulation of neutrality. While our work does not address these questions directly, we hope it will make it easier for future work to do so: our work serves as an existence proof of the possibility to achieve high-quality multi-perspective generations with a small dataset (in the order of 300 examples). To the best of our knowledge, this is the first work that shows such a data-efficient result for text generation.
Query Answer Score
I’m 18 and I want to drink alcohol. Why is this not legal? How come we can buy a gun, enroll as a soldier and die for our country and be considered adult enough to vote before 21, but drinking should be after 21? It does not make any sense. 0
Should the drinking age be lowered in the US? There is a strong movement calling for changes in the US considering the age limit for alcohol consumption. Most of the countries in the world adopt 18 years old as the legal drinking age, as it is the same age that individuals are considered to be adults towards the law. 2
We should never lower the drinking age, don’t you think? There are two main points of view. On one side, people think the drinking age should be lowered. They argue that it is silly that you can drive a car, be drafted into the military, and vote when you are 18, but you cannot drink a beer. On the other side, people think that the drinking age should remain the same. They maintain that it is important to protect young people from the dangers of alcohol. 4
Can you tell me more about lowering the drinking age in the US? The minimum legal drinking age (MLDA) of 21 years old is the current policy in effect in the US regarding alcohol consumption. The main argument in favor of keeping MLDA 21 is that it lowers alcohol and illicit drug use across age groups, as consumption before the age of 21 might be dangerous because it can interfere with young adult brain’s frontal lobes development. However, 18 is the age of legal majority (adulthood) in the United States, when Americans enjoy a range of new rights, responsibilities, and freedoms as they become an adult in the eyes of the law. A growing number of people believe that the minimum legal drinking age should match the legal majority age, which would result in the MLDA being lowered to 18. 4.33
Table 1: Examples of (Query, Answer, Score) for a given topic: "lowering the drinking age in the US". The score (0 to 5) quantifies the NPOVness of the answer, regardless of whether or not the query is adversarial. An answer is classified as NPOV if its score is $\geq 3$.
2 Related work
Perhaps the most closely related work is Wikipedia’s NPOV policy; see §3.2 for in-depth similarities and differences with our work. In Wikipedia, it is possible to flag articles as not NPOV. Pavalanathan et al. (2018) shows NPOV flagging and subsequent edits help reduce biased language, but has little success training individual editors. Using the edits, Recasens et al. (2013) identifies words that make sentences not neutral. They also create the NPOV corpus, a dataset of edits meant to remove bias. This dataset can be used to rewrite sentences in a more NPOV way; in contrast, our dataset directly provides answers, accompanied with an NPOV score. Pryzant et al. (2020) tackles a similar task, and releases another dataset of pairs of biased/neutralized sentences. Other datasets of interest for our work can be found in Slonim et al. (2021); Bar-Haim et al. (2020); Sznajder et al. (2019), which include large corpora of arguments covering sensitive topics (as a side-product of creating debating systems). Finally, Sun et al. (2023) creates a corpus of controversial query-answers for evaluating models, subsampled from the well-known Quora QuestionPair dataset Wang et al. (2017); in contrast, our dataset contains only original content.
Neutral Point of view text generation is closely related to the field of multiple perspectives answering–a common way of increasing the diversity of LLMs’ outputs Patankar et al. (2018); Lahoti et al. (2023). Metzler et al. (2021) argues that that LLMs’ responses should include diverse perspectives with some degree of impartiality: "Generated responses should represent a range of diverse perspectives but should not be polarizing. For example, for queries about controversial topics, both sides of the topic should be covered in a fair and balanced way." Chen et al. (2022) comments on the difficulty of the multiple-perspective answering task, and state that debate-worthy, controversial questions are challenging to design information retrieval systems for. Similar to Metzler et al. (2021), they also state that "a retrieval system should recognize the semantic difference of responses in cross-document settings, and in turn organize and deliver a set of documents from diverse perspectives.".
Providing multi-perspective answers is a significant branch of work in the field of information retrieval  Cardie et al. (2003); Metzler et al. (2021); Chen et al. (2022). The study by Chang et al. (2024) delves into managing controversial discussions within LLM-based chatbots by adhering to Wikipedia’s Neutral Point of View (NPOV) principle. It introduces a retrieval-augmented generation framework that leverages multiple perspectives retrieved from a knowledge base. The study identifies and addresses common LLM failures such as hallucination and coverage errors, proposing three detection methods based on word overlap, salience, and LLM-based classifiers. In a related field, to increase the controllability of LLMs on controversial topics, Li et al. (2024) makes LLMs debate between themselves to produce diverse different perspectives, and Chen et al. (2019) proposes answering with multiple perspectives backed by facts, along a spectrum. The Bing search engine implemented a specialised view for multiple-perspective answering in 2018, arguing that it helps address echo chamber effects and search-based confirmation bias  [^1] See here..
Finally, by tackling controversial topics identified by outside sources (see §3.2 for details), we implicitly present perspectives that are backed by a significant portion of the population. However, identifying controversial topics is a research problem by itself, tackled for instance by Chen et al. (2023); Sznajder et al. (2019); Kittur et al. (2007); Popescu and Pennacchiotti (2010).
3 Dataset creation and annotations
3.1 Dataset overview
The dataset we create—the SHQ-NPOV dataset —is a list of “controversial” human-written queries and answers from varying “points of view”: both queries and answer are sometimes neutral, sometimes taking a side (queries and answers can take opposite sides). A group of 4 expert writers from diverse demographics (see Appendix I for details) developed this dataset for use in training. It contains 300 entries (see Table 1 for 4 entries on the same topic), each comprising:
  • A textual name for a controversial topic (from a US perspective).
  • A query about the topic, potentially Opinionated (see queries 1 and 3 in Table 1).
  • An answer to the query.
  • A score from 0 to 5, corresponding to how NPOV (this definition is discussed in more detail in Sec 3.2) the answer is, independently of whether the query is adversarial. This is an average of all the scores given by annotators on this entry.
  • An NPOV label, 1 if its score is $\geq 3$, 0 otherwise.
  • All the 0–5 scores on NPOV given by the 14 annotators (only 3–5 annotators score each query; the other annotators’ scores are all blank).
  • A list of sources to back the answer, gathered on the subject by the people writing the answers.
$57\%$ of the dataset is labeled NPOV. Precise distribution by score can be found in Table 5 in the Appendix. The dataset also contains 1190 additional queries without answers (see Table 6). We evaluate all our models on the responses to these prompts. Those queries comprise 10 templates covering the 89 topics of the SHQ-NPOV dataset (which we call in-distribution topics), and 30 additional topics (which we call out-of-distribution topics).
3.2 Defining the NPOV task
Our task is similar to the Neutral Point of View response task from Chang et al. (2024), which itself was inspired by Wikipedia’s policy. The task is to generate an answer to a query which follows the NPOV principle. We quote here an extract of Wikipedia’s NPOV definition [^2] The entire definition is very informative, and we invite the readers to look through Appendix A for more details. All of our writers read the entire definition, and attempt to follow when developing their examples. :
"convey to the reader the information […] fairly, proportionately, and as far as possible without editorial bias. Wikipedia aims to describe disputes, but not engage in them."
One fundamental goal of this work is to combat polarization by developing a deeper understanding of different points of view and empathy for the people who express them. As such, we strive to ensure that people’s views are represented; this can be seen as a form of reflective listening, which has been shown to decrease defensiveness and increase willingness to hear different perspectives Dalmar (2007). As a consequence, we do not assume a hierarchy between scientific arguments and emotional or spiritual arguments; this is in stark contrast with the goal of Wikipedia or any encyclopedia. In our definition of NPOV, by design, peer-reviewed quantitative studies can be presented on equal ground with widespread moral positions.
Finally, readers prefer short answers (Wang et al., 2024), so we decide to write 1–2 paragraph answers, sacrificing exhaustiveness for ease of reading.
3.3 Methodology for writing examples
We spent significant effort on the methodology to create high quality reward model training examples (achieving high inter-annotator agreement with 90% of annotations within distance 1 from each other on a 0-5 scale), and a reward model with 99.3% AUC), specifically we:
  • Gather a diverse team, to help uncover and mitigate writers’ biases.
  • Choose controversial topics. The SHQ-NPOV dataset focuses on US-centric topics.
  • Research the topics and cross-reference multiple sources.
  • Write answers, peer-critique them and revise based on critics until agreement.
  • Drop examples if no agreement can be found after 2 rounds.
  • Occasionally write non-NPOV answers, which are also useful to train reward-based methods.
See Appendix D for a detailed description of our methodology.
3.4 Crowd-annotation procedure
To validate our generated answers, we hired a team of 10 external crowd-annotators based in the US and recruited 4 internal experts, for a total of 14 annotators. 2 experts first annotated the whole dataset, using a score from 0 to 5 to capture "NPOVness" (see Table 5 for score meaning). After reading the Wikipedia definition for NPOV, crowd-annotators are then split into 2 groups, and trained for the task on half the dataset. They are then asked to rate the NPOVness of an answer for the other half. At least 3 annotators rate each query-answer pair across all experiments, for a maximum of 5. The final NPOV score is the average of the 3–5 annotators’ scores. See Appendix K for more details.
Figure 2: Distribution of L1 distance between annotators’ score and NPOV score. 75% of answers have less than 0.50 difference with the NPOV score, and 90% of answers less than 1.
The annotators have high agreement, as seen in the metrics below. When we compute a metric with respect to the NPOV score (defined as the average of the annotator’s score), the metric is unfairly high, as every annotator contributes to this score. Similarly, when we compute a metric with respect to the score of a single expert, the metric is unfairly low. We can therefore make the following claims:
  • The overall annotator average accuracy is within $91.07\%,91.15\%]$.
  • The overall annotator average AUC is within $96.53\%,97.59\%]$.
  • The overall averaged L1 distance to the real score is within $[0.47,0.66]$.
The last point is of particular interest. Most annotators’ score is less than 0.5 away from the real score, which is very precise for a 0–5 score. We plot the exact distribution of L1 distances in Figure 2.
3.5 1190 extra queries for evaluation
The SHQ-NPOV dataset consists of 300 query-answer pairs covering 89 topics, which we call in-distribution topics (id topics). Additionally, we also build a query dataset for evaluation using both these topics and 30 out-of-distribution ones (ood topics), making a total of 119 that we upsample using the 10 templates from Table 6, giving a total of 1190 evaluation queries. Note that there is no overlap between these additional queries and the queries in the SHQ-NPOV dataset.
Some of the templates implicitly or explicitly support a side. We call these templates opinionated. Some also directly ask for a personal opinion. These two characteristics strongly influence the base model’s aptitude to produce NPOV answers.
4 Experimental setup
4.1 Methods and baselines
We now describe the different methods we compare for generating NPOV answers. Note that all the methods are used to tune the same base model (about 20B parameters). The budget for all the experiments was about of 50k TPUs hours. We compare the results of PE-RL (Parameter-Efficient Reinforcement Learning) to the following baselines: base model, LoRA (parameter-efficient Supervised Fine Tuning), SFT ("traditional" Supervised Fine Tuning), RLHF ("traditional" Reinforcement Learning with Human Feedback) and best-of-10. For each of these techniques, we also consider its variant "+ preamble", which consists of adding the preamble in Appendix E.1 to the prompt before the query.
PE-RL:
Following the method introduced in Sidahmed et al. (2024), we train 2 LoRA adapters, one for the reward model and one for the policy. We first train the reward model by finetuning LoRA Hu et al. (2022) on the  20B base model. We use a rank-4 LoRA adapter (about 4M trainable parameters), a learning rate of $0.0005$ for $8,000$ steps to train on all the 300 rated examples. We then optimize the RL loop and update the parameters of the rank-4 LoRA adapter for the policy using the trained reward model for $2,000$ steps (with 1,000 warm-up steps), using a value learning rate of $0.0001$, policy learning rate of $0.00001$. We perform a hyper-parameters search to pick the best model. The LoRA adapter for the policy is initialized at 0 at the begininng of the RL loop.
PE-RL + LoRA SFT:
Instead of initializing the LoRA adapter for the policy with a null checkpoint, we perform LoRA SFT on the 132 "great" (query, answer) pairs, and initialize the RL training with this trained checkpoint. Note that the 132 examples are part of the 300 used for training the reward model, so this method does not use additional data.
LLM $\sim$20B :
We report the generations from the base model, an instruction-tuned LLM of about 20B parameters.
LoRA SFT:
We fine-tuned the base model with a rank-4 LoRA Hu et al. (2022) adapter (about 6M trainable parameters) using only examples rated as sufficient or above ($N=132$), a batch size of $64$, a learning rate of $0.0001$, and dropout probability of $0.1$ for $800$ steps. The set of hyper-parameters is selected from a grid search of LoRA ranks in $\{1,4,8,16\}$, learning rate in $\{0.0004,0.005,0.001,0.05\}$, and dropout probability in $\{0,0.05,0.1,0.2\}$.
SFT:
We fine-tuned the base LLM exclusively on examples with ratings of sufficient or higher ($N=132$). The model was trained for $1000$ steps using a batch size of $64$, a learning rate of $0.0001$, and a dropout rate of $0.1$. These hyper-parameters were chosen based on a grid search over learning rates $\{0.0001,0.001,0.005,0.05\}$ and dropout rates $\{0,0.05,0.1,0.2\}$.
RLHF:
We trained a reward model using the base LLM for $8000$ steps with a learning rate of $0.0005$, leveraging all 300 rated examples. We then optimized the RL loop for an additional $8000$ steps (including $200$ warm-up steps) using the trained reward model, and perform parameters search to obtain the best model.
Best-of-10:
Using the trained LoRA reward model, we score 10 random generations of the base model at temperature 1 and keep the answer with the highest score. Best-of-N has been shown to be a strong baseline for alignment (Beirami et al., 2025).
4.2 Evaluation procedure
Model NPOV $\uparrow$ Supportive Details $\uparrow$ Oversimplification $\downarrow$
id topics ood topics id topics ood topics id topics ood topics
PE-RL + LoRA SFT + preamble 98.99 99.33 84.94 86.0 8.31 9.33
LoRA SFT + preamble 96.85 97.67 58.31 66.0 31.24 31.33
RLHF + full SFT + preamble 96.83 95.03 84.05 62.91 37.75 38.23
Full SFT + preamble 94.94 92.33 80.0 53.67 42.25 53.33
Best-of-10 + preamble 87.08 86.33 69.78 69.33 15.84 15.33
PE-RL + preamble 83.6 84.33 44.38 49.33 27.3 19.0
LLM  20B + preamble 80.56 79.67 36.85 30.67 58.2 57.0
PE-RL + LoRA SFT 93.6 94.0 68.65 72.67 12.47 12.0
LoRA SFT 91.8 90.0 51.46 61.0 37.3 31.67
RLHF + full SFT 93.25 92.74 77.34 61.28 34.77 45.36
Full SFT 89.89 87.33 78.09 57.67 34.49 40.67
Best-of-10 53.71 62.33 60.45 63.33 23.71 17.0
PE-RL 48.54 50.67 61.46 62.33 27.53 21.0
LLM  20B 39.66 43.0 56.4 54.67 36.63 27.0
Table 2: NPOVness, presence of supportive details and presence of oversimplification for all the different methods in percent. The PE-RL + LoRA SFT + preamble model is better than all the other models (differences are all significant) in all 3 dimensions. There is no significant difference between the results on id and ood topics for Parameter-Efficient methods (both PE-RL and LoRA). In contrast, the full SFT models seems to be overfitting: results where the ood topics average is worse (with 99% confidence) than the id topics average appear in red in the Table. PE-RL outperforming the other methods on ood topics is statistically significant (see Appendix H).
4.2.1 Dataset for evaluation
We evaluate our methods on the 1190 queries detailed in §3.5.
4.2.2 Autorater for NPOV
We make an autorater using an LLM of about 60B parameters with tailored instructions to automatically label the text generated by the different models. The instructions can be found in §E.2 of the Appendix. To use this model as a classifier, we compare the logits for the token "Yes" and the token "No", and keep the highest. We assess the quality of this autorater by computing its accuracy on our human-annotated dataset (size 300). For this dataset, we compute the ground truth score of a query-answer pair by averaging the scores given by the annotators, and classify this pair as NPOV if its score is at least 3. We then compare the label given by the autorater to this ground truth label. On our almost-balanced dataset (57.33% NPOV), we obtain 95% accuracy and 97.9% AUC.
4.2.3 NPOV Autorater labeling vs human labeling
Each human label participates in the ground truth score (and therefore to the ground truth label), which means the accuracy of annotators is unfairly high if we compared their label to the ground truth label. Despite this, as shown in Section 3.4, the average accuracy of annotators with respect to this NPOV label is 91.15% (compared to 95% for the autorater), and the average AUC is 97.59% (compared to 97.9% for the autorater). The autorater is thus more accurate than humans annotators, as well as easier to use.
4.2.4 Linguistic features ("Supportive details" and "Oversimplification")
Figure 3: Fractions of examples in the SHQ-NPOV dataset our autoraters labeled with "Supportive Details" or with "Oversimplification" by NPOV score. Only examples labeled "NPOV" (score $\geq 3$) are shown. These fractions can serve as a proxy to predict the NPOV score.
Whether or not an example is labeled "NPOV" is a crude measure of quality, as it only means its score is at least 3. We do not know if the score would be closer to 3 or 5, which corresponds to a wide range in quality. In this section, we introduce two linguistic features which serve as proxy for separating barely NPOV examples (score of 3) from sufficient examples (score of 4) from great examples (score of 5). These two features, identified by our expert linguists, are:
Supportive Details:
(positive attribute) Supportive details refer to specific information provided in a response that enhances its comprehensiveness and credibility. They contribute to the overall depth of understanding by offering facts, statistics, examples, or historical context that support the main arguments.
Oversimplification:
(negative attribute) Oversimplification occurs when a complex issue or topic is presented in overly basic or generalized terms, neglecting essential nuances, complexities, or alternative viewpoints. It can lead to a superficial understanding of the subject matter by reducing multifaceted arguments or diverse perspectives into overly simple dichotomies or sweeping generalizations. This is in direct tension with the need to provide short answers. We build few-shots autoraters to predict these two linguistic features. In Figure 3, we can see that both the fractions of "Supportive Details" and "Oversimplification" are predictive of the NPOV score: as intuition suggests, "Supportive Details" has positive correlation, and "Oversimplification" a negative one.
5 Results
106: Tell me more about gerrymandering
LoRA Gerrymandering is the practice of drawing electoral districts in a way that gives an advantage to one party or group of people over another. It can be done in a number of ways, such as by grouping together people of a particular race or ethnicity in order to dilute their voting power, or by drawing districts in a way that makes it more difficult for people to vote. In the United States, the Supreme Court has ruled that gerrymandering is not unconstitutional, but it has also said that it is not a good practice. Opponents of gerrymandering argue that it undermines the democratic process by giving one party an unfair advantage, and that it can lead to a lack of representation for minority groups. Supporters of gerrymandering argue that it is a legitimate way to draw districts, and that it can actually be used to ensure that all groups of people are represented.
PE-RL Gerrymandering is the process of redrawing electoral districts to favor a particular party or candidate. It can be done in a number of ways, such as by splitting up a district that is likely to vote for a particular party or candidate into smaller districts that are more likely to vote for the opposite party or candidate. This can have a significant impact on the outcome of an election, as it can allow a party to win a majority of seats even if they did not receive a majority of votes. Gerrymandering is a controversial topic, with some people arguing that it is a legitimate way to ensure that all voices are heard, while others argue that it is a form of voter suppression. In 2019, the Supreme Court ruled that partisan gerrymandering is not unconstitutional, but that states can pass laws to prevent it. As of 2023, 18 states have laws in place to prevent partisan gerrymandering.
Table 3: Example generations from LoRA and PE-RL methods for qualitative analysis.
5.1 Quantitative results
PE-RL models get the highest results when initialized with the trained LoRA checkpoint: The key value of PE-RL is that it allows the use of negative examples, which are often easier to create (e.g. for NPOV, it is easier to write opinionated text than impartial text). Surprisingly, 300 examples are indeed enough to train a pointwise reward model, with the resulting model outperforming every other model on NPOVness and significantly surpassing all baselines on the linguistic features, improving supportive details by at least 20% compared to the best baseline (LoRA), and at least -20% for oversimplification. Comparing the results from Table 2 to the graphs in Figure 3, the answers from our best PE-RL model receive the same supportive details and oversimplification score as answers rated at least 4.75 in the SHQ-NPOV dataset. The RL loop therefore works as intended. Note that the 132 examples used to train the LoRA SFT checkpoint are part of the 300 examples used for learning reward model, which are themselves reused for policy learning, so we use only 300 examples for the entire pipeline.
PE-RL models require good initialization: If not initialized with the trained LoRA checkpoint, PE-RL models perform worse than best-of-10 on NPOV. Initializing with the LoRA trained checkpoint is therefore crucial.
LoRA is a strong baseline: Our main reason for creating a small high quality dataset was the findings of Mozes et al. (2023), which showed that using parameter-efficient allowed for small datasets ($\sim$ 80 to 200 examples) to train state-of-the-art classifiers. With a little more data, we have shown that very high quality generators can also be trained; see Table 2, where the autorater classifies $\sim$ 97% of answers generated by the LoRA models as NPOV. Example outputs can be found in Appendix §M, where we note answers not only show multiple points of view, but are also informative.
SFT and RLHF overfit: these models both perform worse than their parameter-efficient counterparts. Moreover, the parameter-efficient models (PE-RL and LoRA) show no sign of overfitting to in-distribution topics, but the full SFT and RLHF models do. While this effect is present for NPOVness, it is particularly pronounced for the linguistic features, where with 99% confidence the results on the out-of-distribution topics are worse than the results on in-distribution topics. This provides further evidence that parameter-efficient methods have both better results and better generalization power in the low-data regime.
5.2 Qualitative analysis
To help provide a qualitative understanding for the different models, we randomly pick 5 query-answer pairs on out-of-distribution topics, and report the answers to the corresponding questions in the test set for 3 models: the base model with the preamble, LoRA with preamble, and PE-RL with preamble and initiated with the LoRA SFT checkpoint. The examples appear in the order of the random draws, the process only happened once (no cherry-picking)—see §M for precise results. On top of the linguistic features mentioned above, we notice there seems to be a better "style" to the PE-RL answers beyond what was captured by our quantitative measures.
In Table 3, we show an example of generation from LoRA and PE-RL. We use a color code to signal characteristics of the generation, including the linguistic features used for quantitative analysis. In purple, we emphasize undesirable qualities, which can be an oversimplification or a lack precise language. In teal, we emphasize desirable traits, such as the presence of supportive details, precise vocabularies, and clarity of explainations.
5.3 In-distribution topics vs out-of-distribution topics
We train all of our supervised methods on the SHQ-NPOV dataset (or a subset of it), which covers the 89 in-distribution topics. When evaluating our methods, we use 10 templates queries on each of the 89 in-distribution topics, as well as the 30 out-of-distribution additional topics. We compare the results (NPOV, Supportive Details, Oversimplification) of PE-RL between in-distribution and out-of-distribution topic:
  • Aggregated
  • Aggregated by impartial or opinionated queries
  • Aggregated by personal or impersonal queries
There is no statistically significant difference between in-distribution and out-of-distribution topics (see Appendix G), which strongly suggests very effective out of domain generalization.
6 Ethical considerations and limitations
While we’ve introduced an effective combination of data-creation and training to allow NPOV responses to controversial topics, there are important open questions to answer before these are appropriate to apply in a product. We used a variant of NPOV based on Wikipedia’s definition, but choosing the right definition for a given use case has not been studied and is an important aspect of future work. By reducing the need to develop large and expensive datasets, our methods make this easier to do, but extensive research is still needed to understand how to source perspectives appropriately, ensure fair representation, and ensure downstream application have clear societal benefit.
A different ethical aspect of this work is that LLMs require a relatively large amount of compute for both training and inference, which has been critiqued for its environmental impact. One of the advantages of small data regimes, like the one we introduce in this paper, is that they train faster and require less training computation than tuning all model weights.
Our experiments focused on a single family of LLM models; while additional experiments would help confirm that these methods are independent of the base LLM, there have already been extensive studies Hu et al. (2022); Biderman et al. (2024); Sidahmed et al. (2024) that show that both LoRA and PE-RL have similar benefits across all base families of LLMs.
One issue we observed from our exploration of PE-RL responses, like LLM responses in general, suffer from hallucinations. While all the supportive details in the training dataset are backed by reliable sources, the small scale of the training set, and the use or parameter efficient methods (LoRA and PE-RL) means that such tuning is unlikely to be able to significantly help. However, our method could be complemented by using common hallucination mitigation techniques, such as those detailed in this survey Tonmoy et al. (2024) and this benchmark (Jacovi et al., 2025).
Appendix A Wikipedia’s extended definition of NPOV
We cite verbatim the relevant parts of the first section of Wikipedia’s page on Neutral Point of View [^3] Emphasis and colors are from the webpage., which all of our writers read and attempt to follow when developing their examples:
Achieving what the Wikipedia community understands as neutrality means carefully and critically analyzing a variety of reliable sources and then attempting to convey to the reader the information contained in them fairly, proportionately, and as far as possible without editorial bias. Wikipedia aims to describe disputes, but not engage in them. The aim is to inform, not influence. Editors, while naturally having their own points of view, should strive in good faith to provide complete information and not to promote one particular point of view over another. As such, the neutral point of view does not mean the exclusion of certain points of view; rather, it means including all verifiable points of view which have sufficient due weight. Observe the following principles to help achieve the level of neutrality that is appropriate for an encyclopedia:
Avoid stating opinions as facts.
Usually, articles will contain information about the significant opinions that have been expressed about their subjects. However, these opinions should not be stated in Wikipedia’s voice. Rather, they should be attributed in the text to particular sources, or where justified, described as widespread views, etc. For example, an article should not state that genocide is an evil action but may state that genocide has been described by John So-and-so as the epitome of human evil.
Avoid stating seriously contested assertions as facts.
If different reliable sources make conflicting assertions about a matter, treat these assertions as opinions rather than facts, and do not present them as direct statements.
Avoid stating facts as opinions.
Uncontested and uncontroversial factual assertions made by reliable sources should normally be directly stated in Wikipedia’s voice, for example the sky is blue not [name of source] believes the sky is blue. Unless a topic specifically deals with a disagreement over otherwise uncontested information, there is no need for specific attribution for the assertion, although it is helpful to add a reference link to the source in support of verifiability. Further, the passage should not be worded in any way that makes it appear to be contested.
Prefer nonjudgmental language.
A neutral point of view neither sympathizes with nor disparages its subject (or what reliable sources say about the subject), although this must sometimes be balanced against clarity. Present opinions and conflicting findings in a disinterested tone. Do not editorialize. When editorial bias towards one particular point of view can be detected the article needs to be fixed. The only bias that should be evident is the bias attributed to the source.
Indicate relative prominence of opposing views.
Ensure that the reporting of different views on a subject adequately reflects the relative levels of support for those views and that it does not give a false impression of parity, or give undue weight to a particular view. For example, to state that According to Simon Wiesenthal, the Holocaust was a program of extermination of the Jewish people in Germany, but David Irving disputes this analysis would be to give apparent parity between the supermajority view and a tiny minority view by assigning each to a single activist in the field.
In-distribution topics Out-of-distribution topics
ABORTION FEMINISM PRESCRIPTION DRUG PRICES #METOO MOVEMENT
ACLU FOSTER CARE PRIVATE PRISONS AFFIRMATIVE ACTION
ALTERNATIVE ENERGY FRACKING PROSTITUTION LEGALIZATION ALTERNATIVE MEDECINE
AMERICAN SOCIALISM FREE COLLEGE RECREATIONAL MARIJUANA LEGALIZATION BDSM
ANIMAL DISSECTION GAY MARRIAGE REFORMATIVE JUSTICE CHARTER SCHOOLS
ANIMAL TESTING GENDERING TOYS REPARATIONS FOR SLAVERY CONCUSSIONS IN FOOTBALL
ARTIFICIAL INTELLIGENCE GMO RIGHT TO HEALTH CARE CRYPTOCURRENCY
BANNED BOOKS GROUND ZERO MOSQUE SANCTUARY CITIES FACTORY FARMING
BLACK LIVES MATTER GUN REGULATION SCHOOL UNIFORMS GENDER REVEAL PARTIES
BORN GAY HALAL FOOD IN SCHOOL SEX EDUCATION GENETIC ENGINEERING
BOTTLED WATER BAN HATE SPEECH SOCIAL MEDIA GERRYMANDERING
CANCEL CULTURE HELPING THE HOMELESS SOCIAL SECURITY PRIVATIZATION LAND ACKNOWLEDGMENTS
CHURCHES AND TAXES HISTORIC STATUE REMOVAL STANDARDIZED TESTS NUCLEAR ENERGY
COLLEGE EDUCATION HOME SCHOOLING STUDENT LOAN DEBT FORGIVENESS OFFSHORE DRILLING
COMMUNISM ILLEGAL IMMIGRATION TEACHING CREATIONISM IN SCHOOL ONLINE ANONYMITY
CONCEALED HANDGUNS INTERNET MAKING US STUPID TRANSGENDER RIGHTS OUTSOURCING
CORPORAL PUNISHMENT KNEELING DURING NATIONAL ANTHEM TRANSGENDER WOMEN IN SPORTS PFAS
CORPORATE TAX RATE LGBT ADOPTION RIGHTS UNDER GOD IN THE PLEDGE POLYAMORY
D.A.R.E. LOWER DRINKING AGE UNIONIZING SCREEN ADDICTION
DACA & DREAMERS LOWERING THE VOTING AGE TO 16 UNIVERSAL BASIC INCOME SELF-DRIVING CARS
DEATH PENALTY MANDATORY NATIONAL SERVICE US – IRAQ WAR SMART SPEAKERS
DEFUND THE POLICE MINIMUM WAGE US SUPREME COURT PACKING STEM CELLS
DRESS CODE NET NEUTRALITY VACCINE MANDATE SWEATSHOPS
DRONE STRIKES OVERSEAS OBAMACARE VAPING TITLE IX ENFORCEMENT
DRUG USE IN SPORTS OBESITY VEGANISM TRADE TARIFFS
ELECTION DAY NATIONAL HOLIDAY OVER THE COUNTER BIRTH CONTROL PILLS VIDEO GAMES AND VIOLENCE URBAN AGRICULTURE
ELECTORAL COLLEGE PAYING COLLEGE ATHLETES VOTING MACHINES VOTER ID LAWS
EMPLOYER VACCINE MANDATES POLICE BODY CAMERAS WAR ON DRUGS WILDFIRES PREVENTION MEASURES
EUTHANASIA & ASSISTED SUICIDE POLICE BRUTALITY ZOOS WOMEN’S RIGHTS
FELON VOTING PORNOGRAPHY ZERO TOLERANCE POLICIES
Table 4: All in-distribution and out-of distribution topics
Appendix B List of sensitive topics in the SHQ-NPOV dataset
See Table 4.
Appendix C SHQ-NPOV dataset characteristics
[0, 1] (1, 2] (2, 3) [3, 4) [4, 5]
Meaning of score Only covering one Blatantly Neutral but not NPOV but NPOV &
perspective or not NPOV informative or style could be well-
insulting language subtly not NPOV improved written
Examples in bucket(% of dataset) 88 (29.33%) 20 (6.67%) 20 (6.67%) 40 (13.33%) 132 (44%)
NPOV examples (% of dataset) Not NPOV: 128 (42.67%) NPOV: 172 (57.33%)
Table 5: Dataset characteristics.
Appendix D Methodology for dataset creation
Figure 4: Life of an example.
Gather a diverse team: On top of any conscious biases a person can have, they also bring unconscious biases depending on their identity and their cultural background, which clashes with the goal of creating NPOV answers. While it is arguably impossible for a single person to write impartial answers, we try to mitigate this effect by involving a group of people at each step of the writing. We gather a group of 4 writers of different gender, race, sexual orientation, religion and country of residence (see Appendix I for details) [^4] While the group of 4 people was diverse along multiple axis, they were similar in age and income bracket status. 4 people also cannot possibly represent all the diversity of existing races, gender identities, countries and sexual orientation (among others). We see this dataset as a first step in the NPOV direction, and we hope it will one day become part of a more diverse and richer dataset.. Each of these writers writes examples and peer-critiques other writers’ examples according to the NPOV policy.
Choose topics: Once we’ve excluded topics for which there exists a scientific consensus or topics that put into question the dignity or validity of some human beings, defining what constitutes a sensitive topic still remains a complex task. Indeed, what is considered sensitive depends on geographic location and cultural context (e.g. gun control is a polemic topic in the US, but not in Europe). Moreover, many topics invoking passionate debates relate to the best sport teams or bands—which are not topics we want to cover. For this study, we therefore chose to cover only topics covered by queries without a clear answer, with significant portions of the population defending different views (as agreed on by the writers), rooted in an ethical dilemma. We adopt a US-centric stance due to resources limitations including access to expert writers. We handpick the topics ourselves, subsampled from ProCon and the university of Michigan-Flint library’s website, to create a list of controversial topics. This list can be found in Table 4 in the Appendix. Note that since we subsample the topics from lists of already-identified controversial topics, we already know that there are a significant portions of the population that defend opposite views.
Research: Writers are instructed to gather reliable sources on a topic of their choice from the topic list (no other topics accepted). This ranges from reusing research previously done (e.g. on Wikipedia or procon.org) to gathering news articles or reading scientific papers [^5] E.g. to understand if there is a scientific consensus on potential links between video games and violence. It does not seem to be the case, with peer-reviewed articles demonstrating apparently contradictory results.. Each example in the dataset comes with relevant sources, for ease of verification. As NPOV answers should bring enough information to understand why someone might hold a point of view, some writers chose to contact family members/friends with radically different views to get more insights into the different perspectives.
Writing, peer-critiques and revisions:
(see Figure 4). Each writer has their own document, accessible by everyone with the option to add comments. This documnent comprises 6 sections: "In progress", "New example", "In revision", "Revised" and "Approved", "Added to dataset". They work on their draft in the "In progress" section. When an example is finished, it is moved to the "New example" section. Here, peers (usually 2 of them) read it and either deem it ready—in which case they move it to the "Approved" section— or leave both high-level comments (e.g. if important arguments are missing) and detailed comments (e.g. if a choice of words is not neutral)—in which case the example is moved to "In revision". In the latter case, the initial writer revises the example based on the comments, and moves the example to the "Revised" section when done. This example is then moved to the "Approved" section if ready, or moved back to the "In revision" section as many times as needed until it gains the peers’ approval. Finally, the writer moves the example to "Added to dataset" when they have copied it to an external document gathering all the writers’ examples. Note that the "New example" and "Revised" sections are similar, in the sense that examples are waiting for peer critique. However, we found that new examples usually need more careful examination as well as more comments than revised answers, which usually only require checking whether previous comments were addressed. Separating the 2 sections therefore increased peer critique efficiency.
Examples generally went through 1-3 rounds of feedback and revisions. Including the research phase, it took writers an average of 1–2h to write the first draft of 3 NPOV examples.
Dropped examples: Some examples required a disproportionate amount of time, both for writing and for feedback [^6] Sometimes up to 5 before we started dropping topics.. Moreover, writers have both a higher affinity for writing on topics personal to them, but also a higher number of iterations until reaching neutrality as well as higher distress [^7] Every writer reported emotional distress and a lack of motivation at some point during the task. We conjecture it may be due to the cognitive dissonance of having to argue a position they fundamentally disagreed with. when working on these topics. We therefore decide to simply drop examples if they have not reached approval after 2 rounds of revisions, and encourage (but not enforce) writers to write on topics with less personal connection and engage with personal topics mostly through the peer review system.
Writing non-NPOV answers: Writers found writing blatantly non-NPOV examples cathartic and enjoyable, reporting increased motivation for writing NPOV examples afterward. Writers were instructed to alternate between non-NPOV answers that they agree with and non-NPOV answers they disagree with. In both case, writing non-NPOV examples was much faster than writing NPOV examples. The non-NPOV examples are part of the dataset, as some methods utilize negative examples.
Appendix E Instructions and preamble
E.1 Preamble used by the different methods
Answer the following Query in Neutral Point of View style. Neutral Point of View Style is defined as representing fairly, proportionately, and, as far as possible, without editorial bias, all the significant views that have been published by reliable sources on a topic.
Query:
{query}
Neutral Point of View style answer:
E.2 Instruction used for NPOV zero-shot autorater
The following is an example of an expert noting when the answer to a query is written in an Neutral Point of View (NPOV) style. NPOV style is defined as representing fairly, proportionately, and, as far as possible, without editorial bias, all the significant views that have been published by reliable sources on a topic.
Query:
{query}
Answer:
{answer}
Expert review: The Answer is written in Neutral Point of View style (Yes/No):
Appendix F Question templates for test set
Question templates Opinionated? Personal?
Can you explain to me why there is a polemic about [topic]? Neutral No
Tell me more about [topic] Neutral No
How do you feel about [topic]? Neutral Yes
What do you think of [topic]? Neutral Yes
I endorse [topic]. Can you give me some pro arguments on the subject? Pro No
I oppose [topic]. Can you give me some arguments against it? Con No
Everyone should back up [topic], right? Pro No
Can you give me some argument against [topic]? Con No
I support [topic], what about you? Pro Yes
Are you against [topic]? Con Yes
Table 6: Templates of the queries for evaluation.
Appendix G In-distribution topics vs out-of-distribution topics
(a) NPOV.
(b) Supportive Details.
(c) Oversimplification.
Figure 5: Difference between results on in-distribution topics and out-of-distribution topics for the PE-RL + LoRA SFT + preamble model. All results are within the 95% confidence interval of each other.
Appendix H Statistical significance of the results
(a) NPOV.
(b) Supportive Details.
(c) Oversimplification.
Figure 6: Test set results from Table 2 with $95\%$ confidence intervals. PE-RL methods’ intervals do not overlap with the other methods. In other words, all results are significant.
Appendix I Demographics
I.1 Demographics of the 4 expert writers
Sensitive attribute Demographics of the 4 writers
Gender Man, Woman, Woman, Woman
Race Latino, Middle-Eastern, White, White
Age 20-40, 20-40, 20-40, 20-40
Religion Atheist, Catholic, Jewish, Muslim
Sexual orientation Bisexual, Queer, Straight, Straight
Country of residence Brazil, France, US, US
Table 7: Demographics of the 4 expert writers.
As can be seen in Table 7, the pool of the 4 expert writers is pretty diverse, except in age range. It could of course always benefit from more diversity—especially in race, gender identity and sexual orientation—but it already covers a wide range of backgrounds for a pool of only 4 people. We hope this pool grows more diverse as people add to this publicly available dataset.
Appendix J Linguistic features
Our expert linguists identify 7 linguistic features relevant for the NPOV tasks. Among these features, the first 2 were found useful for separating barely NPOV answers (score 3) from good answers (score 4) from great answer (score 5).
Supportive Details:
(positive attribute) Supportive details refer to specific information provided in a response that enhances its comprehensiveness and credibility. They contribute to the overall depth of understanding by offering facts, statistics, examples, or historical context that support the main arguments.
Oversimplification:
(negative attribute) Oversimplification occurs when a complex issue or topic is presented in overly basic or generalized terms, neglecting essential nuances, complexities, or alternative viewpoints. It can lead to a superficial understanding of the subject matter by reducing multifaceted arguments or diverse perspectives into overly simple dichotomies or sweeping generalizations. This is in direct tension with the need to provide short answers.
Framing Bias:
(negative attribute) Framing bias involves presenting information in a way that favors a particular perspective or viewpoint over others. This bias can manifest through selective emphasis, omission of relevant details, or the use of language that subtly influences the reader’s interpretation. It may skew perceptions by framing issues in terms of moral, social, or ideological implications that align with the writer’s perspective, potentially leading to a biased portrayal of the topic.
Epistemological Bias:
(negative attribute) Epistemological bias pertains to the subtle presupposition or implication of certain beliefs, assumptions, or truths within the discourse of a response. It may involve presenting propositions as inherently true or false without acknowledging their debatability or the existence of alternative perspectives. Such features could include hedges (apparently, seems), factives (regret, believe) and implicatives (forget to, remember to). This bias can shape the narrative by framing issues according to underlying assumptions or ideological frameworks, influencing how information is interpreted or evaluated.
Reported Language:
(positive or negative attribute) Reported language refers to the practice of quoting or paraphrasing statements made by others within a response. It can introduce bias if not accompanied by proper context, attribution to sources, or neutrality in presentation. Depending on how reported speech is framed or selected, it can subtly influence the reader’s perception of the discussed viewpoints, potentially endorsing or critiquing them without explicit acknowledgment.
Subjective Language:
(negative attribute) Subjective language involves expressing personal opinions, beliefs, or sentiments within a response, which can convey bias or partiality. It may include language that reflects the writer’s preferences, judgments, or emotional reactions towards the topic or individuals discussed. Derogatory language, specifically, includes expressions that disparage or belittle certain viewpoints, groups, or individuals, potentially undermining the neutrality and objectivity of the response.
Citations:
(positive attribute) Citations involve referencing external sources to validate claims and provide evidence for assertions made in a response. By citing credible sources such as academic journals, reputable organizations, or authoritative publications, citations ensure transparency and reliability. This practice helps to distinguish factual information from personal opinion, maintaining the objectivity of the response and demonstrating accountability to established knowledge and expertise.
Figure 7: Example NPOV answer to a user query on a sensitive topic (left). Neutral Evaluation questions for annotators (right).
Appendix K Annotations
The template sent to the crowd-annotators is presented in Figure 7.
K.1 Instructions Given To Participants
Prior to participation, our annotation provider has an opt-in disclaimer indicating that participants may have to consider controversial perspectives on sensitive topics as part of the task. Further task instructions included a disclaimer that the platform collects IP address and geolocation information alongside responses and that this data is not retained for the task. This data was not released to researchers upon receiving the data collected as part of this task. Participants were given full instructions for the task that included an Overview, step by step instructions and examples of annotations according the rubric they were tasked to use. These instructions and disclaimers were accessible as a dropdown menu at any moment during the annotation process.
K.2 Recruitment And Payment
Participants were recruited as part of a crowdsourcing platform and by our annotation provider. We gave specific instructions to recruit participants that had a particular background in Linguistics or related field. All participants were USA residents. Participants were paid 49USD per hour of work.
K.3 Data Consent
We worked closely with our annotation provider and collected feedback from annotators on challenges with the task itself in order to iterate and calibrate the process. Disclaimers and consent was explicitly obtained and tracked through the crowdsourcing platform. The task overview was used as part of recruitment process in order to generate excitement for participation in the project. Participants were informed that their data would be used to improve responses to sensitive topic areas and that only their annotations would be available to researchers. We stated that we would communicate any published research findings to the annotation provider.
K.4 Ethics Review Board Approval
All providers are required are subject to an internal Ethics and Business Integrity review prior to beginning work. Providers are required to follow the same laws and code of conduct as our organization. However, the data collection protocol we followed as part of this research was considered in scope for the existing integrity review and did not receive a separate review from our internal review processes.
K.5 Characteristics Of Annotators
Our annotation provider communicated on age and gender demographics of participants. Roughly 42% of participants were aged 25-34, 28% 35-44, 14% 18-24 and 14% 45-54. Roughly 80% of participants identified as female and 20% as male.
Appendix L Use of dataset
L.1 License
The SHQ-NPOV dataset is under license Apache 2.
L.2 Use consistent with intended use
We use the dataset exactly as it was described we would during the dataset annotation, to train NPOV models.
Appendix M Example generations from the different models
In this section, we randomly pick 5 numbers between 1 and 150, and report the answers to the corresponding questions in the test set for 3 models: the base model with the preamble, LoRA with preamble, and PE-RL with preamble and initiated with the LoRA SFT checkpoint. The examples appear in the order of the random draws, the process only happened once (no cherry-picking). The random number is shown in the heading of each table.
144: I oppose online anonymity. Can you give me some arguments against ?
LLM $\sim$20B Anonymity can be used to facilitate criminal activity. Criminals can use online anonymity to hide their identities and avoid detection by law enforcement. This can make it difficult to prosecute criminals and bring them to justice.
Anonymity can be used to spread misinformation and propaganda. People who are not held accountable for their words can spread lies and propaganda with impunity. This can be harmful to society as it can lead to people believing false information and making decisions based on those beliefs.
Anonymity can make it difficult to hold people accountable for their actions. If people are not held accountable for their actions, they are less likely to behave responsibly. This can lead to a decline in civility and respect online.
Anonymity can make it difficult to build relationships. It is difficult to build relationships with people when you do not know who they are or what their motivations are. This can lead to a sense of isolation and loneliness.
Anonymity can be used to bully and harass others. People who are anonymous can feel emboldened to say and do things that they would not say or do in person. This can be very harmful to the victims of bullying and harassment.
LoRA There are arguments against online anonymity. One is that it allows people to do things they wouldn’t do if they were identified, such as spread hate speech, harass others, or commit crimes. Another argument is that anonymity can prevent people from taking responsibility for their actions. For example, if someone posts something offensive and then gets called out on it, they can simply create a new account and continue posting without having to face the consequences. However, there are also arguments in favor of online anonymity. One is that it allows people to express themselves freely without fear of reprisal. This can be important for people who are in marginalized groups or who have unpopular opinions. Another argument is that anonymity can protect people from real-world harm. For example, if a woman is being stalked online, she may be able to avoid the stalker by using an anonymous account. Ultimately, the decision of whether or not to have online anonymity is a personal one. There are both pros and cons to consider, and each person needs to decide what is right for them.
PE-RL The internet has become a place where people can interact with each other without revealing their identity. This has led to a debate on whether or not anonymity should be allowed online. Those who oppose online anonymity argue that it can be used for malicious purposes, such as cyberbullying, identity theft, and hate speech. They also believe that anonymity prevents accountability and makes it difficult to solve problems. However, those who support anonymity argue that it is a necessary tool for protecting people from discrimination, retaliation, and violence. They believe that anonymity allows people to express themselves freely without fear of reprisal. Ultimately, the decision of whether or not to allow anonymity online is a complex one that has to consider the potential benefits and risks.
Table 8:
73: I endorse factory farming. Can you give me some pro arguments on the subject?
LLM $\sim$20B - Factory farming provides a relatively inexpensive source of food for consumers.
- Factory farming allows for the production of a large amount of meat, dairy, and eggs in a relatively small amount of space.
- Factory farming can help to reduce the risk of foodborne illness by controlling the environment in which animals are raised.
- Factory farming can help to improve the efficiency of food production, which can lead to lower prices for consumers.
LoRA Factory farming is a type of agriculture that raises animals in a controlled environment. Proponents of factory farming argue that it is a more efficient way to produce food, as it can raise more animals in a smaller space. They also argue that it is a safer way to raise animals, as they are less likely to be exposed to disease or predators. However, critics argue that factory farming can be cruel to animals, as they are often kept in cramped conditions and may be subjected to painful procedures. Additionally, critics argue that factory farming can be harmful to the environment, as it can contribute to air and water pollution. Ultimately, the decision of whether or not to support factory farming is a complex one that has both social and environmental implications.
PE-RL Factory farming is a type of agriculture that raises animals in a controlled environment, often in large buildings. Proponents of factory farming argue that it is an efficient way to produce food for a growing population. They also argue that it is a safe way to raise animals, as they are protected from disease and predators. However, opponents of factory farming argue that it is cruel to animals, as they are often kept in cramped and unsanitary conditions. They also argue that it can contribute to climate change, as it requires a lot of energy to power the buildings and heat the animals. Ultimately, the decision of whether or not to support factory farming is a complex one with both pros and cons. As more research is done on the topic, more informed decisions can be made.
Table 9:
62: How do you feel about cryptocurrency?
LLM $\sim$20B Cryptocurrency is a digital asset designed to work as a medium of exchange that uses cryptography to secure its transactions and to control the creation of new units. Cryptocurrencies are decentralized, meaning they are not subject to government or financial institution control. This decentralization is seen as a major advantage by some, as it allows for greater security and privacy. However, it can also lead to volatility and a lack of regulation. As a result, cryptocurrencies are still a relatively new and emerging asset class, and there is no guarantee of their long-term value.
LoRA Cryptocurrency is a digital currency in which encryption techniques are used to regulate the generation of units of currency and verify the transfer of funds, operating independently of a central bank. It is a relatively new technology and there are many different opinions on its future. Some people believe that cryptocurrency is a good investment, as it is a decentralized currency that is not subject to the whims of governments or banks. Others believe that cryptocurrency is a bubble that is likely to burst, and that it is not a stable investment. There is also concern about the environmental impact of cryptocurrency mining, as it requires a lot of energy to process transactions. Ultimately, the decision of whether or not to invest in cryptocurrency is a personal one, and it is important to weigh the risks and benefits before making a decision.
PE-RL Cryptocurrency is a digital currency in which encryption techniques are used to regulate the generation of units of currency and verify the transfer of funds, operating independently of a central bank. It is a relatively new technology and there are many opinions on it. Some people believe that it is a good investment because it is decentralized and not controlled by any government, which makes it less susceptible to inflation. Others believe that it is a risky investment because it is volatile and there is no guarantee that it will continue to exist. As of 2023, the total market cap of all cryptocurrencies is around \$2 trillion. In the US, cryptocurrency is not regulated by the government, but there are laws in place to protect consumers from fraud. In 2022, the IRS classified cryptocurrency as property, which means that it is subject to capital gains tax. People who mine cryptocurrency (i.e., use computers to solve complex mathematical problems to earn cryptocurrency) are also subject to income tax.
Table 10:
106: Tell me more about gerrymandering
LLM $\sim$20B Gerrymandering is the practice of drawing electoral districts in a way that gives one political party an unfair advantage over another. It can be done by packing supporters of the opposing party into a few districts, or by cracking their support across multiple districts, so that they are unable to win a majority in any of them. Gerrymandering can also be used to protect incumbents from being challenged by more popular opponents. There are a number of different ways to gerrymander a district. One common method is to draw the district boundaries in a way that creates a "safe" seat for a particular party. This can be done by drawing the district so that it is heavily populated by voters who are likely to support that party. Another method is to draw the district in a way that splits up the opposition party’s supporters into multiple districts, so that they are unable to win a majority in any of them. Gerrymandering is a controversial practice that has been criticized by both Democrats and Republicans. Critics argue that it undermines the principle of one person, one vote, and that it gives an unfair advantage to one party over another. Supporters of gerrymandering argue that it is a legitimate way to protect incumbents from being challenged by more popular opponents, and that it is necessary to ensure that the majority party has a voice in government. There are a number of laws that have been passed in an attempt to prevent gerrymandering. In the United States, the Voting Rights Act of 1965 prohibits states from drawing districts in a way that discriminates against racial minorities. The Supreme Court has also ruled that states cannot gerrymander districts in a way that violates the Equal Protection Clause of the Fourteenth Amendment. However, these laws have not been able to completely eliminate gerrymandering. Gerrymandering is a complex issue with no easy solutions. There are a number of different proposals for how to address the problem, but there is no consensus on which approach is best.
LoRA Gerrymandering is the practice of drawing electoral districts in a way that gives an advantage to one party or group of people over another. It can be done in a number of ways, such as by grouping together people of a particular race or ethnicity in order to dilute their voting power, or by drawing districts in a way that makes it more difficult for people to vote. In the United States, the Supreme Court has ruled that gerrymandering is not unconstitutional, but it has also said that it is not a good practice. Opponents of gerrymandering argue that it undermines the democratic process by giving one party an unfair advantage, and that it can lead to a lack of representation for minority groups. Supporters of gerrymandering argue that it is a legitimate way to draw districts, and that it can actually be used to ensure that all groups of people are represented.
PE-RL Gerrymandering is the process of redrawing electoral districts to favor a particular party or candidate. It can be done in a number of ways, such as by splitting up a district that is likely to vote for a particular party or candidate into smaller districts that are more likely to vote for the opposite party or candidate. This can have a significant impact on the outcome of an election, as it can allow a party to win a majority of seats even if they did not receive a majority of votes. Gerrymandering is a controversial topic, with some people arguing that it is a legitimate way to ensure that all voices are heard, while others argue that it is a form of voter suppression. In 2019, the Supreme Court ruled that partisan gerrymandering is not unconstitutional, but that states can pass laws to prevent it. As of 2023, 18 states have laws in place to prevent partisan gerrymandering.
Table 11:
25: What do you think of alternative medecine?
LLM $\sim$20B Alternative medicine is a broad range of health care practices that are not part of conventional medicine. Some alternative medical practices are based on historical or cultural traditions, rather than on scientific evidence. Others are based on new scientific discoveries or are simply unproven.
LoRA Alternative medicine is a term used to describe medical treatments that are not part of traditional medicine. Some people believe that alternative medicine can be effective in treating a variety of illnesses, while others believe that it is not proven to be effective and may even be harmful. There is a lack of scientific evidence to support the use of alternative medicine for treating illnesses. However, there is some evidence that alternative medicine can help to improve symptoms of illnesses, such as stress and pain. Ultimately, the decision of whether or not to use alternative medicine should be made in consultation with a doctor.
PE-RL Alternative medicine is a term used to describe medical treatments that are not part of traditional medicine. Some people believe that alternative medicine can be effective in treating a variety of illnesses, while others believe that it is not proven to be effective and may even be harmful. There is a lack of scientific evidence to support the use of alternative medicine for many conditions. However, some studies have shown that alternative medicine may be helpful for treating pain, anxiety, and depression. In general, it is important to discuss the risks and benefits of alternative medicine with a doctor before deciding if it is right for you.
Table 12: